OpenStudy (anonymous):

integral (pi/4 to pi/2) of cos(x)/sin(x)

OpenStudy (anonymous):

\[\int\limits_{\pi/4}^{\pi/2}\frac{ \cos(x) }{ \sin(x) }\]

OpenStudy (amistre64):

since cos is a derivative of sine ....

OpenStudy (anonymous):

and i know you use substitution and you insert u=sin(x) and d(u)=cos(x) and it turns into \[\int\limits_{?}^{?}\frac{ 1 }{ u }\] and then it turns into \[\ln \left| u \right|\] and then it turns into \[\ln \sin(x) from \frac{ \pi }{ 4 } \to \frac{ \pi }{ 2}\]

OpenStudy (anonymous):

but you plug it in and you get \[0-\ln \frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

but i think this is wrong

hartnn (hartnn):

its correct, it just needs simplification.

OpenStudy (anonymous):

how?

hartnn (hartnn):

what is \(\large \dfrac{2}{\sqrt2}=...?\) can you simplify this ? and write it in the form 2^(.....)

OpenStudy (anonymous):

You can also use the property of logarithm that \[\ln{\frac{a}{b}}=\ln a -\ln b\]

OpenStudy (anonymous):

and \[\ln a^b = b\ln a\] as well

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } \ln 2 - \ln 2\]

hartnn (hartnn):

the log property that you'll need is \(\large \log a^b=b\log a\)

OpenStudy (anonymous):

They are fundamentally the same.

OpenStudy (amistre64):

if we keep it in u u = sinx ; x=pi/4 .... u = sqrt(2)/2 u = sinx; x=pi/2 .... u = 1 1 - ln(sqrt(2)/2) is correct

OpenStudy (anonymous):

i dont really understand

hartnn (hartnn):

did you forget - *minus*

OpenStudy (amistre64):

i dont think it needs to be "simplified" any. the calculator knows how to work it as is :)

hartnn (hartnn):

ln 1 =0

OpenStudy (amistre64):

.... i meant that lol

OpenStudy (anonymous):

its not in the multiple choice

OpenStudy (anonymous):

\[-\ln (\frac{\sqrt 2}{2})=-\ln(\frac{1}{\sqrt{2}})=-\ln(2)^{-\frac{1}{2}}=\frac{\ln 2}{2}\]

OpenStudy (amistre64):

whats in the options then?

OpenStudy (amistre64):

we can manipulate this thing all day, but it might go quicker if we knew the options

hartnn (hartnn):

that was negative sign before also \(-(\frac{ 1 }{ 2 } \ln 2 - \ln 2)= \ln 2- \frac{ 1 }{ 2 } \ln 2 = \frac{ 1 }{ 2 } \ln 2\)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

how

hartnn (hartnn):

\(\frac{ 1 }{ 2 } \ln 2 = \ln 2^{\frac{ 1 }{ 2 } }=\ln \sqrt2\)

hartnn (hartnn):

i used the same log property \(\large \log a^b=b\log a\)

hartnn (hartnn):

did you get it ?

OpenStudy (anonymous):

yup thanks

hartnn (hartnn):

welcome ^_^